# Dissipationless dynamics of spin supersolid states in a spin-1/2 triangular antiferromagnet with impurities

### v1

@@ Abstract
Motivated by recent experimental observations of possible spin supersolid states in triangular lattice compounds, we study the dynamical properties of various ground states in the spin-1/2 easy-axis antiferromagnetic Heisenberg model with impurities under magnetic fields, using numerical Density Matrix Renormalization Group methods. For both spin supersolid states in the low and high fields, the gapless Goldstone mode at the $K$ points remains robust with impurities, which is related to the presence of spin superfluidity. As a comparison, we find a splitting of magnon band at the same impurity density level in the conventional magnetic state, the so-called up-up-down state. In addition, the finite superfluid stiffness probed by the twisted phase in the spin supersolid states is consistent with the excitation spectrum. We argue that this excitation spectrum with impurity provides direct evidence of the dissipationless dynamics in the spin supersolid states, which could be tested in neutron scattering experiments.
@@
Introduction.— The supersolid features coexisting superfluidity and spatial symmetry breaking order which is originally proposed as an exotic quantum state in Helium [1, 2, 3, 4, 5]. Recently, distinctive manifestation of the supersolidity has also been discovered in the ultracold quantum gases, resulting in a dipolar supersolid [6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16]. Given the fact that the boson models can be mapped onto the spin models, the spin supersolid may exist in frustrated spin systems, with the triangular lattice Heisenberg antiferromagnets as the most promising platform [17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28]. To this end, previous numerical studies [29, 30, 31, 32] have shown spin supersolid phases in the weak and strong magnetic field regime, and an up-up-down (UUD) phase between them. A magnetic field-induced phase diagram has been mapped out both at zero and finite temperatures [29, 32].
@@
Figure 1: Panel (a) shows the triangular lattice with blue solid circles represent positive spin values, and the red shaded circles represent negative spin values. The purple solid circles represent the impurity sites with a positive spin value. The radius represents the magnitude where the purple ones have $\left\langle S_{i}^{z}\right\rangle\approx 0.5$. The ground state is obtained at $h_{z}=0.836$ on the $N=48\times 6$ lattice where only the middle part is shown. Panel (b) shows the illustration of a triangular lattice. Panel (c) shows the corresponding Brillouin zone and momentum cuts for the dynamical structure factor.
@@
The triangular compound $\text{Na}_{2}\text{BaCo}(\text{PO}_{4})_{2}$, which was extensively studied as a candidate for quantum spin liquids [33, 34, 35, 36, 37, 38], has attracted further interest recently because of the potential realization of spin supersolid states under magnetic fields [32, 39, 40, 41, 42, 43, 44, 45, 46, 47]. In particular, a giant magnetocaloric effect is observed in the supersolid regime [40] which promotes high-performance demagnetization cooling [44, 48], and the phase diagram agrees well with the numerical studies of the spin-1/2 anisotropic triangular Heisenberg model. Thus, this material provides an ideal platform to explore the spin supersolid states. Further exciting progress from inelastic neutron scattering experiments show the low-energy excitations with rotonlike minimum at the M points and gapless Goldstone modes at the K points [41, 47], both of which are consistent with the dynamical spin structure factor calculated with unbiased numerical approaches [31, 41, 47]. However, such rotonlike minimum can also exist in quantum spin liquids [49, 50] that the spin supersolid state might be close to [51]. Despite extensive efforts, direct experimental observation of the superfluidity in the spin supersolid states remains an open question.
@@
One of the key characteristics of superfluidity is the dissipationless dynamics that results from the scattering of the spin supercurrent [52, 53]. Indeed, recent spin current studies through the spin Seebeck effect have revealed a saturating supercurrent at low temperatures [54, 55] and a long-distance transport of the thermally induced spin current [54]. However, the incoherent magnons might also be injected into the system through the thermal methods that lead to condensation. On the other hand, the scattering due to spin supercurrent is independent of local impurities, and the low-energy excitations in the dynamical spin structure factor would remain the same in the presence of impurities. In particular, the robust Goldstone mode at the $K$ points with impurities is directly related to the superfluid density. In contrast, for the UUD state the impurities could drastically change the low-energy spectrum.
@@
Motivated by the recent experimental realization of spin supersolids, we numerically study the spin-1/2 anisotropic triangular Heisenberg model with magnetic fields. We show consistent results of the superfluid stiffness in the supersolid phases at both zero temperature and finite temperature, which could guide experimental searches for the signals of dissipationless dynamics due to spin superfluidity. Most importantly, we propose that the dissipationless dynamics could be identified by the dynamical spin structure factor which shows the robust gapless Goldstone mode even with impurities, which is in sharp contrast to the UUD phase where the lower bands split at the $K$ points with the same impurity density. The dynamical spin structure factor may be measured in the inelastic neutron scattering experiments where the impurities are introduced through element substitution.
@@
Model and methods.— We study the nearest-neighbor spin-1/2 antiferromagnetic Heisenberg model on a triangular lattice, where the Hamiltonian is defined as
@@
Here $\left\langle ij\right\rangle$ refers to the nearest neighbor sites and $J=1$ as the energy unit. To be applicable to the compound $\text{Na}_{2}\text{BaCo}(\text{PO}_{4})_{2}$, we set $\Delta_{z}/J=1.68$ which is determined in Ref. [32] by fitting the experimental data of magnetic specific heat and magnetic susceptibility.
@@
The time evolution is implemented using the time-dependent variational principle (TDVP) for both real and imaginary time [60, 61, 62]. For the zero-temperature real-time dynamics, we employ the one-site TDVP scheme with an enlarged bond dimension achieved via global Krylov vectors [63]. $D=2200$ are used to simulate the time up to $\tau_{tot}=50/J$.
For finite-temperature calculations, we employ imaginary-time evolution using thermal tensor network [62, 64] to construct the density matrix $\rho(\beta)\equiv e^{-\beta H}$ [65]. Simulations were performed on $N=18\times 6$ cylinders. We retain $D=2000$ bond states, implement U(1) symmetry, and achieve a truncation error of $\epsilon\lesssim 5\times 10^{-5}$. The bond dimension is enlarged through the controlled bond expansion algorithm [66, 67].
Superfluid density.— The magnetic field-induced phase diagram at zero temperature has been carried out in previous studies [29, 32, 39, 40] for the easy-axis triangular antiferromagnetic Heisenberg model showing a “Y” supersolid state, an UUD state, a “V” supersolid state, and a polarized state. These states are characterized by $\left\langle m_{z}^{2}\right\rangle$ and $\left\langle m_{\bot}^{2}\right\rangle$ [22], which are related to the Bragg peaks of the spin structure factor at $K$ points via $\left\langle m_{z}^{2}\right\rangle=S^{z}(\textbf{K})/L_{y}^{2}=\frac{1}{L_{y}^{4}}\sum_{i,j}e^{i\textbf{K}\cdot(\textbf{r}_{i}-\textbf{r}_{j}))}\left\langle S^{z}_{i}S^{z}_{j}\right\rangle$ and $\left\langle m_{\bot}^{2}\right\rangle=S^{xy}(\textbf{K})/L_{y}^{2}=\frac{1}{L_{y}^{4}}\sum_{i,j}e^{i\textbf{K}\cdot(\textbf{r}_{i}-\textbf{r}_{j})}\left\langle S^{x}_{i}S^{x}_{j}+S^{y}_{i}S^{y}_{j}\right\rangle$, where the summation is over the middle $L_{y}\times L_{y}$ sites. As shown in Figs. 2(a) and (b), $\left\langle m_{z}^{2}\right\rangle$ and $\left\langle m_{\bot}^{2}\right\rangle$ are finite in the supersolid states, and simultaneously reach maximum and minimum in the UUD states between $B_{1c}\approx 1.49$ and $B_{2c}\approx 4.15$, respectively. Above $B_{3c}\approx 6.54$ the state becomes polarized. This is consistent with previous work. The finite $\left\langle m_{\bot}^{2}\right\rangle$ in the UUD state results from the quantum fluctuations and it becomes smaller on the wider $L_{y}=9$ lattice; see SM [56] for more details.
@@
where $F(\theta)$ is the free energy for a given $\theta$ defined as $F=-\frac{1}{\beta}\log{Z}$ where $Z={\text{Tr}}[\rho(\beta/2)\rho^{\dagger}(\beta/2)]$, and at zero temperature it is equivalent to the ground state energy $E_{0}(\theta)$. We choose $\theta=\pi$ where the energy difference is orders of magnitude larger than numerical accuracy; see more discussion in the SM [56]. As shown in Fig. 2(c), for zero temperature the $\Delta F(\pi)=\Delta E_{0}(\pi)$ increases with the magnetic field in the “Y” supersolid phase, and decreases to zero as the phase approaches the UUD. A peak value of $\Delta E_{0}(\pi)$ is obtained around $h_{z}/J=0.836$ in the “Y” state. At higher fields, the $\Delta E_{0}(\pi)$ becomes finite in the “V” supersolid state with a peak around $h_{z}/J=4.82$ before vanishing for $h_{z}/J>6.54$ in the polarized state. A small kink is identified around $h_{z}/J=5.5$ in both $\Delta E_{0}(\pi)$ and $\left\langle m_{\bot}^{2}\right\rangle$ which is also observed on wider $L_{y}=9$ lattices; see more details in SM [56].
To examine whether the superfluid density remains finite in experimentally accessible temperatures, we obtain the $\Delta F(\pi)$ at finite temperatures. As shown in Fig. 2(c), at low temperatures the evolution of $\Delta F(\pi)$ is qualitatively the same as the one at zero temperature, and the $\Delta F(\pi)$ becomes much smaller for higher temperatures. For both “Y” and “V” states, the maximum value of $\Delta F(\pi)$ appears at the same $h_{z}$ for both zero and finite temperatures. In the finite temperature results of Fig. 2(d), the domes of finite $\Delta F(\pi)$ indicate the “Y” and “V” states, which is consistent with the classical picture [32]. The two domes are separated by the UUD phase where $\Delta F(\pi)$ remains zero within the numerical accuracy. The $\Delta F(\pi)$ persists up to $T/J\approx 0.1$ in the spin supersolid phase; see more details in SM [56].
@@
To provide further theoretical understanding of the dynamical spin structure factor, we obtain linear spin wave results that can capture the low energy excitation dispersions and the isolated magnon branches. Without impurity, the semiclassical ground state contains a three-site unit cell, where the classical spin configuration is obtained by minimizing the energy before performing the Holstein-Primakoff transformation; see more details in SM [56].
Figures 3(a) and (c) show the $\chi(\textbf{q},\omega)$ near the $K$ points obtained in the “Y” supersolid and “V” supersolid phase, respectively. The paths in the Brillouin zone are illustrated in Fig. 1(c). The supersolid states exhibit gapless Goldstone modes from spontaneous U(1) symmetry breaking at the $K$ points with high concentration of the spectral weight while the UUD state only has gapped spin excitations, which is consistent with previous results [74, 75, 49, 31]. This is the key feature related to the superfluidity, and is consistent with the lowest magnon branch from the linear spin wave theory. In the presence of finite impurities, the gapless mode at the $K$ points remains robust, as shown in Figs. 3(d) and (f) for the “Y” and “V” states, respectively. This is in contrast to usual gapless states where the lowest energy excitations are affected the most by disorder or impurity, and it indicates the dissipationless dynamics as an intrinsic property associated with the spin supersolid states. Interestingly, we observe a small gapped mode in the “V” state near the $K$ points, as shown in Fig. 3(c). This may relate to the pseudo-Goldstone mode that results from the three-fold degeneracy of the diagonal order via the order-by-quantum-disorder mechanism [76]. The three-fold degeneracy refers to $\uparrow\uparrow\downarrow$, $\uparrow\downarrow\uparrow$, and $\downarrow\uparrow\uparrow$ in a three-site unit cell. This pseudo-Goldstone mode has also been observed in the zero-field “Y” state [41] with a smaller gap, where the ground state has six-fold degeneracy. However, it is not shown in the “Y” state in Fig. 3(a) which may result from the finite numerical resolution of the spectrum. With impurities we find that the gapped mode near the $K$ points in the “V” state disappears in Fig. 3(f), as the impurities break the degeneracy of the diagonal order.
As a comparison, we obtain the spectrum in the UUD state. As shown in Fig. 3(b), there is no continuum excitation spectrum in the UUD state. Because all spins align in the $z$ direction in the ground state, the magnons are excited in the transverse plane which is mostly captured by the linear spin wave theory, though a renormalization of the magnon dispersion is observed at higher energy which may be due to the interactions between the single quasiparticles. In the presence of impurities, the lowest two magnon bands split near the $K$ points as shown in Fig. 3(e).
@@
To further study the impurity effect of high symmetry points we obtain the $\chi(\textbf{q},\omega)$ near the $M$ points. As shown in Fig. 4(a), in the “Y” supersolid state the rotonlike minimum at the $M$ points is found to be almost flat in the spectral weight, which is consistent with previous study using the infinite projected entangled-pair state (iPEPS) methods [31]. We notice that the flatness is not observed in the linear spin wave dispersion, and is caused by the interactions between the single magnon branches. However, the excitation energy at the $M$ points is closer to the linear spin wave results in this easy-axis Heisenberg model, as compared to the near isotropic one [77, 75]. With impurities, the spectral weight at the $M$ points broaden as shown in Fig. 4(d), which may decay into higher energy modes due to impurity. Similar broadening effect due to the impurity can be found in the “V” supersolid state, as shown in Figs. 4(c) and (f). Except for the broadening effect, the $\chi(\textbf{q},\omega)$ remains almost the same in the presence of impurity. On the other hand, a band splitting is identified at the lowest energy in the UUD state with impurity, which can be seen by comparing Figs. 4(b) and (d).
Summary.— Through extensive numerical simulations on width-6 cylinders, we have explored the dynamical spin structure factors in various magnetic field-induced ground states in the easy-axis triangular-lattice Heisenberg antiferromagnets. In particular, we develop a numerical characterization of the dissipationless dynamics in the spin supersolid state through the excitation spectrum with impurities. We show a robust gapless Goldstone mode at the $K$ points in the presence of finite impurities in the spin supersolid phase, which provides direct evidence of the superfluidity that can be observed in neutron scattering experiments. As a comparison, for the UUD state we find that the lower bands split with the same impurity density. For the spectral weight at higher energies such as the rotonlike minimum, the impurities cause a broadening effect, while the overall profile of the spectrum remains almost unchanged. We believe that our methods could also be applied to other spin supersolid candidate materials, such as $\text{K}_{2}\text{Co}(\text{SeO}_{3})_{2}$ [78, 79, 80, 81, 82, 83] and $\text{Na}_{2}\text{BaNi}(\text{PO}_{4})_{2}$ [84, 85].
@@
For finite-temperature calculations, the procedure starts from a high-temperature expansion of the density matrix $\rho(\beta_{0})\simeq 1-\beta_{0}H+\frac{\beta_{0}^{2}}{2}H^{2}$ [65, 62], with $\beta_{0}=2^{-15}$. We then successively double the inverse temperature until $\beta=1$, followed by a linear evolution in $\beta$ with step size $\delta\beta=1$ down to the lowest temperature. In practice, we employ the one-site tangent space tensor renormalization group scheme with U(1) symmetry on the $N=18\times 6$ lattice with bond dimensions up to $D=2000$. The bond dimension is enlarged through the controlled bond expansion algorithm [66, 67], with an increment of $\delta D=200$ and a truncation error maintained at $\epsilon\lesssim 5\times 10^{-5}$.
The dynamical spin structure factor is calculated from the time dependent spin correlations as given by the Eq. 4 in the main text. Due to limited simulation time, we apply a smearing factor of $e^{-\eta\tau}$ to the time series before performing the discrete Fourier transformation where $\eta=1/\tau_{tot}$ and $\tau_{tot}$ are the total simulation time. To evaluate the convergence of the dynamical spin structure factor, we compare the simulations with different maximum bond dimensions. For a direct comparison we use the same $\tau_{tot}$ for the calculation with different maximum bond dimensions. Comparing Figs. S1 (a) and (c) with Figs. S1(b) and (d) as an example, the dynamical spin structure factor with bond dimensions up to $D=2200$ is almost the same as the one with up to $D=1400$. This is mainly because large bond dimensions are only needed at the later time where the entanglement entropy increases, while the convolution with the smearing factor has distributed the main weight to the simulation data at the early time.
@@
The impurities are placed evenly in the bulk of the lattice as shown in Figs. S2(a), (b), and (c) for the “Y” supersolid state, the up-up-down (UUD) state, and the “V” supersolid state, respectively. The four impurities in the summation of the bulk $N_{mid}=\frac{3}{4}L_{x}\times L_{y}$ sites correspond to an impurity density level of $1.85\%$. For simplicity, we only consider the impurity on the same sublattice where the same magnetic structure in $z$ direction is pinned by the impurities. In this case, the $\left\langle m_{z}^{2}\right\rangle$ remains almost the same. For a general impurity configuration where impurities are placed randomly on the three different sublattices, the interference of impurities may reduce the $\left\langle m_{z}^{2}\right\rangle$ but the impact to superfluid stiffness remains similar [73].

### v2

@@ Abstract
Motivated by recent experimental observations of the possible spin supersolid states in triangular lattice compounds, we study the dynamical properties of various ground states in the spin-1/2 easy-axis antiferromagnetic Heisenberg model with impurities under magnetic fields using numerical methods. In both low- and high-field spin supersolid states, the gapless Goldstone mode at the $K$ points remains robust against impurities, which is related to the presence of spin superfluidity. By contrast, we find that impurities induce a splitting of the magnon bands at the same density level in the conventional magnetic state, the so-called up-up-down state. In addition, the finite superfluid stiffness probed by the twisted phase in the spin supersolid states is consistent with the excitation spectrum. We argue that this excitation spectrum with impurity provides direct evidence for the dissipationless dynamics in the spin supersolid states, which could be tested in neutron scattering experiments.
@@
Introduction.— The supersolid features coexisting superfluidity and spatial symmetry breaking order which is originally proposed as an exotic quantum state in solid Helium [1, 2, 3, 4, 5]. Recently, distinctive manifestation of the supersolidity has also been discovered in the ultracold quantum gases, resulting in the dipolar supersolid [6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16]. Given the fact that the boson models can be mapped onto the spin models, the spin supersolid may exist in frustrated spin systems, with the triangular lattice Heisenberg antiferromagnets as the most promising platform [17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28]. To this end, previous numerical studies [29, 30, 31, 32] have shown spin supersolid states in the weak and strong magnetic field regimes, and an up-up-down (UUD) state between them. The phase diagram by tuning magnetic fields has been mapped out both at zero and finite temperatures [29, 32].
@@
Figure 1: Panel (a) shows the triangular lattice with blue solid circles representing positive spin values, and red shaded circles representing negative spin values. The purple solid circles represent the impurity sites with a positive spin value. The radius represents the magnitude where the purple ones have $\left\langle S_{i}^{z}\right\rangle\approx 0.5$. The ground state is obtained at $h_{z}=0.836$ on the $N=48\times 6$ lattice where only the middle part is shown. Panel (b) shows the illustration of a triangular lattice. Panel (c) shows the corresponding Brillouin zone and momentum cuts for the dynamical spin structure factor.
@@
The triangular compound $\text{Na}_{2}\text{BaCo}(\text{PO}_{4})_{2}$, which was extensively studied as a candidate for quantum spin liquids [33, 34, 35, 36, 37, 38], has attracted further interest recently because of the potential realization of spin supersolid states under magnetic fields [32, 39, 40, 41, 42, 43, 44, 45, 46, 47]. In particular, a giant magnetocaloric effect is observed in the critical regime [40] which promotes high-performance demagnetization cooling [44, 48], and the phase diagram agrees well with the numerical studies of the spin-1/2 anisotropic triangular Heisenberg model. Thus, this material provides an ideal platform to explore the spin supersolid states. Further exciting progress from inelastic neutron scattering experiments shows the low-energy excitations with rotonlike minimum at the M points and gapless Goldstone modes at the K points [41, 47], both of which are consistent with the dynamical spin structure factor calculated with unbiased numerical approaches [31, 41, 47]. However, such rotonlike minimum can also exist in quantum spin liquids [49, 50] that the spin supersolid state might be close to [51]. Despite extensive efforts, direct experimental observation of the superfluidity in the spin supersolid states remains an open question.
@@
One of the key characteristics of superfluidity is the dissipationless dynamics that results from the scattering of the spin supercurrent [52, 53]. Indeed, recent spin current studies through the spin Seebeck effect have revealed a saturating supercurrent at low temperatures [54, 55] and a long-distance transport of the thermally induced spin current [54]. However, the incoherent magnons might also be injected into the system through the thermal methods that lead to condensation. On the other hand, the scattering due to spin supercurrent is independent of local impurities, and the low-energy excitations in the dynamical spin structure factor would remain the same in the presence of impurities. In particular, the robust Goldstone mode at the $K$ points against impurities is directly related to the superfluid density. By contrast, for the UUD state the impurities could drastically change the low-energy spectrum.
@@
Motivated by the recent experimental realization of spin supersolids, we numerically study the spin-1/2 anisotropic triangular Heisenberg model with magnetic fields. We show consistent results of the superfluid stiffness in the supersolid phases at both zero temperature and finite temperature, which could guide experimental searches for the signals of dissipationless dynamics due to spin superfluidity. Most importantly, we propose that the dissipationless dynamics could be identified by the dynamical spin structure factor which shows the robust gapless Goldstone mode even with impurities, which is in sharp contrast to the UUD phase where impurities induce a splitting of the magnon bands at the $K$ points at the same density level. The dynamical spin structure factor may be measured in inelastic neutron scattering experiments where the impurities are introduced through the element substitution.
@@
Model and methods.— We study the spin-1/2 easy-axis antiferromagnetic Heisenberg model on a triangular lattice, where the Hamiltonian is defined as
@@
Here $\left\langle ij\right\rangle$ refers to the nearest neighbor sites and $J$ is set to 1 as the energy unit. To be applicable to the compound $\text{Na}_{2}\text{BaCo}(\text{PO}_{4})_{2}$, we set $\Delta_{z}/J=1.68$ which is determined in Ref. [32] by fitting the experimental data of magnetic specific heat and magnetic susceptibility.
@@
The time evolution is implemented using the time-dependent variational principle (TDVP) for both real and imaginary time [60, 61, 62]. For real time dynamics of ground states, we employ the one-site TDVP scheme with an enlarged bond dimension achieved via global Krylov vectors [63]. $D=2200$ are used to simulate the time up to $\tau_{tot}=50/J$.
For finite temperature calculations, we employ imaginary time evolution using thermal tensor network [62, 64] to construct the density matrix $\rho(\beta)\equiv e^{-\beta H}$ [65]. Simulations were performed on $N=18\times 6$ cylinders. We retain $D=2000$ bond states, implement U(1) symmetry, and achieve a truncation error of $\epsilon\lesssim 5\times 10^{-5}$. The bond dimension is enlarged through the controlled bond expansion algorithm [66, 67].
Superfluid stiffness.— The phase diagram for finite magnetic fields at zero temperature has been carried out in previous studies [29, 32, 39, 40] for the easy-axis triangular antiferromagnetic Heisenberg model showing a “Y” supersolid state, an UUD state, a “V” supersolid state, and a polarized state. These states are characterized by $\left\langle m_{z}^{2}\right\rangle$ and $\left\langle m_{\bot}^{2}\right\rangle$ [22], which are related to the Bragg peaks of the spin structure factor at $K$ points via $\left\langle m_{z}^{2}\right\rangle=S^{z}(\textbf{K})/L_{y}^{2}=\frac{1}{L_{y}^{4}}\sum_{i,j}e^{i\textbf{K}\cdot(\textbf{r}_{i}-\textbf{r}_{j}))}\left\langle S^{z}_{i}S^{z}_{j}\right\rangle$ and $\left\langle m_{\bot}^{2}\right\rangle=S^{xy}(\textbf{K})/L_{y}^{2}=\frac{1}{L_{y}^{4}}\sum_{i,j}e^{i\textbf{K}\cdot(\textbf{r}_{i}-\textbf{r}_{j})}\left\langle S^{x}_{i}S^{x}_{j}+S^{y}_{i}S^{y}_{j}\right\rangle$, where the summation is over the middle $L_{y}\times L_{y}$ sites. As shown in Figs. 2(a) and (b), $\left\langle m_{z}^{2}\right\rangle$ and $\left\langle m_{\bot}^{2}\right\rangle$ are finite in the supersolid states, and simultaneously reach maximum and minimum in the UUD states between $B_{1c}\approx 1.49$ and $B_{2c}\approx 4.15$, respectively. Above $B_{3c}\approx 6.54$ the state becomes polarized. This is consistent with previous work. The finite $\left\langle m_{\bot}^{2}\right\rangle$ in the UUD state results from the quantum fluctuations and it becomes smaller on the wider $L_{y}=9$ lattice; see SM [56] for more details.
@@
where $F(\theta)$ is the free energy for a given $\theta$ defined as $F=-\frac{1}{\beta}\log{Z}$ where $Z={\text{Tr}}[\rho(\beta/2)\rho^{\dagger}(\beta/2)]$, and at zero temperature it is equivalent to the ground state energy $E_{0}(\theta)$. We choose $\theta=\pi$ where the energy difference is orders of magnitude larger than numerical accuracy; see more discussion in SM [56]. As shown in Fig. 2(c), for zero temperature the $\Delta F(\pi)=\Delta E_{0}(\pi)$ increases with the magnetic field in the “Y” supersolid phase, and decreases to zero as the phase approaches the UUD. A peak value of $\Delta E_{0}(\pi)$ is obtained around $h_{z}/J=0.836$ in the “Y” state. At higher fields, the $\Delta E_{0}(\pi)$ becomes finite in the “V” supersolid state with a peak around $h_{z}/J=4.82$ before vanishing for $h_{z}/J>6.54$ in the polarized state. A small kink is identified around $h_{z}/J=5.5$ in both $\Delta E_{0}(\pi)$ and $\left\langle m_{\bot}^{2}\right\rangle$ which is also observed on wider $L_{y}=9$ lattices; see more details in SM [56].
To examine whether the superfluid density remains finite in experimentally accessible temperatures, we obtain the $\Delta F(\pi)$ at finite temperatures. As shown in Fig. 2(c), at low temperatures the evolution of $\Delta F(\pi)$ is qualitatively the same as the one at zero temperature, and the $\Delta F(\pi)$ becomes much smaller for higher temperatures. For both “Y” and “V” states, the maximum value of $\Delta F(\pi)$ appears at the same $h_{z}$ for both zero and finite temperatures. In the finite temperature results of Fig. 2(d), the domes of finite $\Delta F(\pi)$ indicate the “Y” and “V” states, which is consistent with the classical picture [32]. The two domes are separated by the UUD phase where $\Delta F(\pi)$ remains zero within the numerical accuracy. The $\Delta F(\pi)$ persists up to $T/J\approx 0.1$ in the spin supersolid states; see more details in SM [56].
@@
To provide further theoretical understanding of the dynamical spin structure factor, we obtain linear spin wave theory results that can capture the low energy excitations and the isolated magnon branches. Without impurity, the semiclassical ground state contains a three-site unit cell, where the classical spin configuration is obtained by minimizing the energy before performing the Holstein-Primakoff transformation [74]; see more details in SM [56].
Figures 3(a) and (c) show the $\chi(\textbf{q},\omega)$ near the $K$ points obtained in the “Y” supersolid and “V” supersolid state, respectively. The paths in the Brillouin zone are illustrated in Fig. 1(c). The supersolid states exhibit gapless Goldstone modes from spontaneous U(1) symmetry breaking at the $K$ points with high concentration of the spectral weight while the UUD state only has gapped spin excitations, which is consistent with previous results [75, 76, 49, 31]. This is the key feature related to the superfluidity, and is consistent with the lowest magnon branch from the linear spin wave theory. In the presence of finite impurities, the gapless mode at the $K$ points remains robust, as shown in Figs. 3(d) and (f) for the “Y” and “V” states, respectively. This is in contrast to usual gapless states where the lowest energy excitations are affected the most by disorder or impurity, and it indicates the dissipationless dynamics as an intrinsic property associated with the spin supersolid states. Interestingly, we observe a small gapped mode in the “V” state near the $K$ points, as shown in Fig. 3(c). This may relate to the pseudo-Goldstone mode that results from the three-fold degeneracy of the diagonal order via the order-by-quantum-disorder mechanism [77]. The three-fold degeneracy refers to $\uparrow\uparrow\downarrow$, $\uparrow\downarrow\uparrow$, and $\downarrow\uparrow\uparrow$ in a three-site unit cell. This pseudo-Goldstone mode has also been observed in the zero-field “Y” state [41] with a smaller gap, where the ground state has six-fold degeneracy. However, it is not shown in the “Y” state in Fig. 3(a) which may result from the finite numerical resolution of the spectrum. With impurities we find that the gapped mode near the $K$ points in the “V” state disappears in Fig. 3(f), as the impurities break the degeneracy of the diagonal order.
As a comparison, we obtain the spectrum in the UUD state. As shown in Fig. 3(b), there is no continuum excitation spectrum in the UUD state. Because all spins align in the $z$ direction in the ground state, the magnons are excited in the transverse plane which is mostly captured by the linear spin wave theory, though a renormalization of the magnon dispersion is observed at higher energy which may be due to the interactions between the single quasiparticles. With the same impurity density level, the lowest two magnon bands split near the $K$ points as shown in Fig. 3(e).
@@
To further study the impurity effect of high symmetry points we obtain the $\chi(\textbf{q},\omega)$ near the $M$ points. As shown in Fig. 4(a), in the “Y” supersolid state the rotonlike minimum at the $M$ points is found to be almost flat in the spectral weight, which is consistent with previous study using the infinite projected entangled-pair state (iPEPS) methods [31]. The flatness is not observed in the linear spin wave dispersion, and is caused by the interactions between the single magnon branches. However, the excitation energy at the $M$ points is closer to the linear spin wave results in this easy-axis Heisenberg model, as compared to the near isotropic one [78, 76]. With impurities, the spectral weight at the $M$ points broaden as shown in Fig. 4(d), which may decay into higher energy modes due to the impurity. Similar broadening effect induced by the impurity can be found in the “V” supersolid state, as shown in Figs. 4(c) and (f). Except for the broadening effect, the $\chi(\textbf{q},\omega)$ remains almost unchanged in the presence of impurity. On the other hand, a band splitting is identified at the lowest energy in the UUD state with impurity, which can be seen by comparing Figs. 4(b) and (d).
Summary.— Through extensive numerical simulations on width-6 cylinders, we have investigated the dynamical spin structure factors in various magnetic field-induced states in the easy-axis triangular-lattice Heisenberg antiferromagnets. In particular, we develop a numerical characterization of the dissipationless dynamics in the spin supersolid state through the excitation spectrum with impurities. We show a robust gapless Goldstone mode at the $K$ points against finite impurities in the spin supersolid states, which provides direct evidence for the spin superfluidity that can be observed in neutron scattering experiments. By contrast, for the UUD state we find that the lower bands split at the same impurity density. For the spectral weight at higher energies such as the rotonlike minimum, the impurities cause a broadening effect, while the overall profile of the spectrum remains almost unchanged. We believe that our results could also be applied to other spin supersolid candidate materials, such as $\text{K}_{2}\text{Co}(\text{SeO}_{3})_{2}$ [79, 80, 81, 82, 83, 84] and $\text{Na}_{2}\text{BaNi}(\text{PO}_{4})_{2}$ [85, 86, 87] as long as the gapless Goldstone mode can be observed.
@@
For finite temperature calculations, the procedure starts from a high-temperature expansion of the density matrix $\rho(\beta_{0})\simeq 1-\beta_{0}H+\frac{\beta_{0}^{2}}{2}H^{2}$ [65, 62], with $\beta_{0}=2^{-15}$. We then successively double the inverse temperature until $\beta=1$, followed by a linear evolution in $\beta$ with step size $\delta\beta=1$ down to the lowest temperature. In practice, we employ the one-site tangent space tensor renormalization group scheme with U(1) symmetry on the $N=18\times 6$ lattice with bond dimensions up to $D=2000$. The bond dimension is enlarged through the controlled bond expansion algorithm [66, 67], with an increment of $\delta D=200$ and a truncation error maintained at $\epsilon\lesssim 5\times 10^{-5}$.
The dynamical spin structure factor is calculated from the time dependent spin correlations as given by the Eq. 4 in the main text. Due to limited simulation time, we apply a smearing factor of $e^{-\eta\tau}$ to the time series before performing the discrete Fourier transformation where $\eta=1/\tau_{tot}$ and $\tau_{tot}$ is the total simulation time. To evaluate the convergence of the dynamical spin structure factor, we compare the simulations with different maximum bond dimensions. For a direct comparison we use the same $\tau_{tot}$ for the calculation with different maximum bond dimensions. Comparing Figs. S1 (a) and (c) with Figs. S1(b) and (d) as an example, the dynamical spin structure factor with bond dimensions up to $D=2200$ is almost the same as the one with up to $D=1400$. This is mainly because large bond dimensions are only needed at the later time where the entanglement entropy increases, while the convolution with the smearing factor has distributed the main weight to the simulation data at the early time.
@@
The impurities are placed evenly in the bulk of the lattice as shown in Figs. S2(a), (b), and (c) for the “Y” supersolid state, the up-up-down (UUD) state, and the “V” supersolid state, respectively. The four impurities in the summation of the bulk $N_{mid}=\frac{3}{4}L_{x}\times L_{y}$ sites correspond to an impurity density level of $1.85\%$. For simplicity, we only consider the impurity on the same sublattice where the same magnetic structure in $z$ direction is pinned by the impurities. In this case, the $\left\langle m_{z}^{2}\right\rangle$ remains almost unchanged. For a general impurity configuration where impurities are placed randomly on the three different sublattices, the interference of impurities may reduce the $\left\langle m_{z}^{2}\right\rangle$ but the impact to superfluid stiffness remains similar [73].